MICROWAVE SPECTROSCOPY OF THE NaCl-(H_{2}O)_{n} (n = 1, 2, 3) COMPLEXES

\underline{ASAO MIZOGUCHI} and YASUKI ENDO, Department of Pure and Applied Sciences, Graduate School of Arts and Sciences, The University of Tokyo, Tokyo 153-8902 Japan; YA-SUHIRO OHSHIMA, Department of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502 Japan.

We report a microwave spectroscopic study of the NaCl-(H_{2}O)_{n} (n = 1, 2, 3) complexes. In the frequency region from 5 to 25 GHz, rotational spectra of NaCl-(H_{2}O)_{n} (n = 1, 2, 3) have been observed for the first time by using a Fourier-transform microwave spectrometer coupled to a laser ablation source. Molecular constants including the nuclear quadrupole coupling constants of Na and Cl nuclei were precisely determined for each complex. We observed various isotopic species of each complex to determine the molecular structures. NaCl-H_{2}O has a nearly planar and cyclic (Na-Cl-H-O) structure, NaCl- (H_{2}O)_{2} has a C_{2} structure with two equivalent H_{2}O, and NaCl-(H_{2}O)_{3} is a symmetric top molecule with three equivalent H_{2}O. The determined structures clearly show that the bond length r(NaCl) increases monotonically with the number of H_{2}O. The determined nuclear quadrupole coupling constnats of the Na and Cl nuclei show considerable changes of charge distribution around these atoms.